T т, D >LOAD

In the class E chopper ( Prof ) the equation of current I minimum and I maximium using any mathematical (Laplace, integration and defferential)
Answer #1

LOAD D2 Ty T2 Calculation Iminimum and I maximum 8- Fos the Load , nohich mna be DC Tnotor e= Ri+ L di +E O St tON them nehanFron Etion = Vo-E L- In 73+)S R+S.L Vs- E S R S L Laplace nverse dor aluove Vc-E 1- e O tTon the dime doman Exeb reeiion fo sat t To T- TON t=Imm - --TON)/Ta Imn- -(T-ToN)/Ta ) hchere Equation cam be salved lor Ime amd and vi ToN/TO 3-3 TON/TO Imx =Subslilutong Imex rem qnaton -(T-Ton) To Vs R - (T-)E TON/To -E 1- e 1-eTO/To ToN Ta ToN ta 4 R To